OpenStudy (anonymous):

Perform the operation and simplify. (z + 5)(z - 7)

OpenStudy (anonymous):

Using the distributive property, you will multiply z by z and -7 and you will multiply 5 by z and -7. z(z) + z(-7) + 5(z) + 5(-7) = z^2 + -7z + 5z + -35 Now, combine the like terms, which are both z's. -7z + 5z = -2z So, your answer will be z^2 - 2z -35.